[2x+9]+7=3 has 2 solutions because of the constant on the right side of the equations is positive or is it 1 solutions please help me explain
Answer by Alan3354(69443)

[2x+9]+7=3 has 2 solutions because of the constant on the right side of the equations is positive or is it 1 solutions please help me explain
------------
[2x+9]+7=3
2x+9 = -4
2x = -13
x = -13/2 one solution.